### v2

# Tur[k]ingBench: A Challenge Benchmark for Web Agents

Abstract
Recent chatbots have demonstrated impressive ability to understand and communicate in raw-text form. However, there is more to the world than raw text. For example, humans spend long hours of their time on web pages, where text is intertwined with other modalities and tasks are accomplished in the form of various complex interactions. Can state-of-the-art multi-modal models generalize to such complex domains?
To address this question, we introduce TurkingBench, a benchmark of tasks formulated as web pages containing textual instructions with multi-modal context. Unlike existing work which employs artificially synthesized web pages, here we use natural HTML pages that were originally designed for crowdsourcing workers for various annotation purposes. The HTML instructions of each task are also instantiated with various values (obtained from the crowdsourcing tasks) to form new instances of the task. This benchmark contains 32.2K instances distributed across 158 tasks.
Additionally, to facilitate the evaluation on TurkingBench, we develop an evaluation framework that connects the responses of chatbots to modifications on web pages (modifying a text box, checking a radio, etc.). We evaluate the performance of state-of-the-art models, including language-only, vision-only, and layout-only models, and their combinations, on this benchmark. Our findings reveal that these models perform significantly better than random chance, yet considerable room exists for improvement. We hope this benchmark will help facilitate the evaluation and development of web-based agents.111Dataset and baselines are available at: https://turkingbench.github.io
1 Introduction
There has been tremendous progress in building diverse AI models, thanks to the availability of large pre-trained language models (LLM) (Radford et al., 2019). However, the bulk of the progress has been limited to chatbots that process raw text-form in their input or output. However, humans rarely work with text-only interfaces. Instead, they generally like to work with visually situated language, such as web-based applications where language is intertwined with other visual elements such as images and tables. It is only recently that multi-modal chatbots have entered the arena (OpenAI, 2023; Anil et al., 2023; Kil et al., 2024), and plenty is yet to be explored in terms of their capabilities and limits, particularly in the context of web-based exploration. A key limitation is the lack of well-defined and natural benchmarks that encompass the complex, interactive, and multi-modal reasoning for various web-based tasks that provide an objective measure of progress for web-navigating agents.
In this work, we introduce TurkingBench, a benchmark of 158 web-grounded tasks. Examples of these tasks are shown in Figure 1. The information content in these tasks is relayed via various stylistic elements (colors, font sizes, shapes, etc.), structural knowledge (tables, paragraph headings, etc.), as well as modalities (text, images, videos, etc.) to communicate various complex tasks for human users. A thorough understanding of these multi-modal signals is a complex task that requires various skills, including understanding language situated in rich visual cues.
The tasks in TurkingBench are collected from tasks published on the Amazon Mechanical Turk (AMT) crowdsourcing platform, a well-established platform for connecting human workers to solve micro-tasks in return for a reward. We are inspired by the ability humans to solve a variety of tasks through web-based interfaces (Efrat & Levy, 2020).
There are also other distinguishing factors to TurkingBench in comparison to the existing literature. The instructions for solving our tasks are interleaved within each web page, and solving them require effective understanding of the page content. This is distinct from most related benchmark (Deng et al., 2023; Zhou et al., 2023) where the task instructions are provided as as a free-form sentences. Regarding the naturalness of the data, the web-pages in TurkingBench were originally tailored for human crowd workers, making them more natural compared to the recent benchmarks that engineer simplified web pages (Liu et al., 2018; Furuta et al., 2023). Last but not least, since the tasks were originally intended for solving non-trivial annotation tasks on crowdsoucing platforms (e.g., building complex NLP benchmarks) they inherently posses a high degree of complexity. We discuss more related work in §2.
There are a variety of challenges involved in solving our proposed task: (i) Multi-modality—these tasks require the models to “see” or “hear” the task content. (ii) Interaction— The agents’ actions are executed in interaction with the web pages, taking into account the context of the previous changes. (iii) Long context—the web pages (the HTML encoding, screenshots, etc.) are typically longer than the context window of existing models.
We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1). We observe that GPT4 can perform notably higher than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page. However, GPT4’s performance is notably far from the estimated upper bound of the dataset. We also analyze model performance as a function of field type and task length, which identify challenges to our future progress.
In summary, we introduce TurkingBench, a benchmark of web-based tasks that requires multi-modal and interactive reasoning on web pages. We formalize the evaluation framework for interacting with and modifying these pages. The introduction of this benchmark aims to encourage the development of agent models that are not only capable of understanding language within a multi-modal context, but also able to transform their knowledge into a sequence of affordances when interacting with web pages. We evaluate notable models on this benchmark and show that they have noteworthy room for progress. We hope this benchmark motivates and enables further development of assistive, web-based agents.
| MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2023) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
| Inputs | language commands | language commands | customer service queries | shopping queries | language commands | English commands | instructions embedded in web pages |
| The input construction | crowdsourced | manual | mined online | crowdsourced | author-written & LM-augmented | author-written | collected from MTurk |
| Environments | simplified web page | simplified but compositional | real-world websites | shopping web pages | real-world websites | variety of web pages | crowdsourcing websites |
| Natural inputs? | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ |
| Realistic interface? | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ | ✓ |
| Dynamic interaction | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
| Functional correctness? | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
2 Related Work
Web-navigation benchmarks.
There are several benchmarks for quantifying abilities of web-navigation agents (Lù et al., 2024; Koh et al., 2024; He et al., 2024). Table Table 1 shows a comparison among a subset of these works. These benchmarks provide an easy way to evaluate models in a reproducible manner. However, all these benchmarks are limited compared to the real-world web. For example, most of the existing datasets operate on websites with limited complexity, such as the ones that are pre-specified, synthesized, or simplified (Liu et al., 2018; Li et al., 2020; Yao et al., 2022; Furuta et al., 2023; Cheng et al., 2023). In almost all the existing benchmarks, the task definitions (instructions) are done after collecting the websites, through manual crowdsourcing or written manually. However, in TurkingBench, instructions are also naturally mined, as they are already embedded in the crowdsourcing web pages. It is also worth noting a limitations of TurkingBench compared to other benchmarks: here the tasks are limited to one page (no page navigation). Given these trade-offs, the field is likely to benefit from all the aforementioned benchmarks.
Instruction-following datasets.
Being able to execute actions on the web also connects to instruction-following literature. Currently, the dominant majority of this literature has has focused on raw text instructions and outputs (Wang et al., 2022; Mishra et al., 2022; Bach et al., 2022; Dwivedi-Yu et al., 2022; Ouyang et al., 2022; Sanh et al., 2022; Shen et al., 2023). The problem of web-exploration (and the proposed benchmark here) is a strict extension of this literature with multiple modalities, with large inputs (long HTML or large screenshots) and strategic decision-making.
Multi-modal reasoning tasks.
TurkingBench is also related to the efforts on multi-modal reasoning. Some works focus on grounding commands to web page elements, setting interaction aside (Pasupat et al., 2018). Several recent efforts study the problem in interactive settings (Gur et al., 2018; Ku et al., 2020; Li et al., 2020; 2022; Li & Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021), though their instructions are simple, typically only a few sentences long. Our work differs since our instructions are part of the web pages, and inferring them involves multi-modal reasoning.
Web-based agents.
The idea of intelligent automated assistant agents that collaborates with humans to complete tasks has been around (Allen et al., 2007). One can view web-based agents as extensions of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022). However, general-purpose web-browsing necessitates dealing with challenges well beyond the classic semantic parsing; for example, the diversity of web pages (no two pages are identical), their change over time, the vast action space (a variety of ways that one can interact with web), multi-modality (text, vision, audio), interactive nature of web-browsing, and so on.
The existing literature use a variety of supervision protocols for supervision, for example, behavior cloning (supervised learning of actions) (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022), in-context prompting and demonstrations (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023). Here are a few highlighted works, among others: (Sodhi et al., 2023) proposed web agents equipped with hierarchical policies (HeaP), i.e., low-level actions that are guided by high-level plans. (Liu et al., 2023) benchmarks a variety of open-source and open-source and closed-source (API-based) models on several well-known interactive benchmarks. While the focus of our work is introducing a new benchmark, we have intended to provide a diverse variety of models to better understand its difficulty. Our efforts are not exhaustive and leave more room for future work to explore more effective model engineering.
3 TurkingBench: Benchmarking Web Agents with Multi-Modal, Interactive Turking Tasks
We propose the development of TurkingBench. The first purpose of this benchmark is to serve as a standardized evaluation of model generalizability. In addition to supporting evaluation, a subset of this data can streamline model design and development.
3.1 Benchmark Schema
TurkingBench consists of a collection of tasks where each task is a bundle of the following components:
A web template containing instructions for the task, input variables, and input fields for the outputs.
Input values to be instantiated for the input variables.
Annotated output labels provided by crowd workers.
An example is shown is shown in Figure 1. As the example shows, the HTML template contains variables (e.g., ${sys10}) that are instantiated with input values. Furthermore, the HTML template contains input fields for receiving input values from the web page users. For each field, we have values previously collected from crowd workers that we will use for evaluation.
3.2 Collecting the tasks
For a benchmark of tasks grounded in web pages to effectively measure progress and generalizability, it needs to be diverse and with broad coverage. Toward such a recourse, we plan on utilizing two existing sources: (A) We will collect the crowdsourcing templates and the resulting annotations from various AI researchers who are active requesters on AMT. (B) Another incidental resource is the set of tasks published on AMT sandbox,222https://requestersandbox.mturk.com a testing environment for requesters to test out the prototypes of their crowdsourcing templates. Since the sandbox environment contains tasks from all users of AMT, it offers more variety than (A). However, they would require additional work to annotate and clean up.
Ultimately, we put together a collection of 158 crowdsourcing tasks (examples in Figure 1). These include but are not limited to language processing tasks (such as paraphrasing, validating factuality, entailment, sentiments, classification, dialogue quality, and rationale generation) and tasks that involve processing images or videos.
3.3 Statistics
Figure 3 shows the overall statistics of TurkingBench. The dataset contain a large sum of input-output instances (36.2K) across 158 tasks. Furthermore, Figure 2 shows the distribution of the input fields represented in our benchmark. The most common field is “radio,” which is expected as it is commonly used to identify preferences. Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
| Measure | Value |
| # of tasks | 158 |
| # of instances | 36.2K |
| avg. # of fields per task | 15.6 |
| avg. length (subwords) of the tasks | 16.8K |
3.4 Programmatic Interaction with Web
We developed a Python library (Figure 4) to streamline the interaction of self-supervised models with our tasks and their evaluations (Figure 4). Our library contains various components. To serve our tasks, we use Turkle333https://github.com/hltcoe/turkle which is an open-source replication of AMT. The content of these tasks are then accessed through the web-browsers that are loaded by Selenium.444https://github.com/SeleniumHQ/selenium
Furthermore, we have developed a library of actions for accessing the pages’ content and making changes to them. This is discussed in §3.4.1. The starting point of interaction is our evaluation script that loops over the tasks and their instances. This is fleshed out in §3.6
3.4.1 A Library of Web Actions
To support model design, we have developed a library of “actions” that can execute various operations on a web page. Our set of actions wrap around the API libraries provided by Selenium (c.f. footnote 4) and PyAutoGUI555https://github.com/asweigart/pyautogui these are sophisticated libraries for web manipulation and the amount of sophistication these libraries provide is significantly more than what is here. Therefore, we build our wrappers around these actions to build an action library with enough complexity to cover the dominant majority of tasks in our benchmark.
| Actions | Modality | Description |
| modify_text | T | modifies the text of input box |
| modify_checkbox | T | modifies the selection of checkbox |
| modify_radio | T | modifies a radio button |
| modify_select | T | selects an item in a drop-down menu |
| modify_range | T | modifies a range input |
| get_html | T | fetches the HTML content of a page |
| capture_screen | V | fetches the screenshot of a page |
| click | V | clicks on a given coordinate |
| type | V | keyboard type at the selected |
| scroll | – | scrolls up or down |
| maximize | – | maximizes the web page |
The list of the actions included in our library is shown in Table 3. In terms of the modalities of information, a subset of these actions are either concerned with executing tasks in text (HTML) modality or visual modality. There are fewer actions in the visual modality since visually many actions are combinations of clicking and typing. This is unlike the text modality where various element types require slightly different treatment.
In terms the nature of their roles, one can split these actions into three groups: (i) Modification actions allow the models to modify each page’s input elements. (ii) Navigation actions allow the models to explore the page, for example, by scrolling up/down or waiting for all elements to be loaded. (iii) Sensing actions allow models to retrieve the latest information about the web page, for example, by fetching its HTML code, taking screenshots of the active page, or around a given element.
Web interaction as “tool” resolution.
The overall task defined here can be viewed as iterative tool resolution based on the context of each task. A web agent solving this task must use a sequence of actions from our action library, based on the context of the task’s instruction (see an example in Figure 5). This connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2023; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023) that use Tools to ground LM generations in physical environments. The set of actions defined in our action library is the set of “tools” that can be executed to obtain specific outcomes. The core of this setup is built upon an LM aware of the tools, the problem context, and the ultimate goal.
3.5 Evaluation Metrics
We devise an evaluation metric that is sensitive to the output type:
Text fields: For a given text response, we compute ROUGE metric (Lin, 2004) against each label and compute their maximum value.
Radio/select fields: We compute an exact match of the predicted label against the majority vote of gold labels.
Checkbox fields: The results of the checkbox selection is a set. Therefore, we quantify this as intersection over union metric between the two sets (set of predicted labels and the gold labels).
Range fields: The range predictions provide real or integer values. We compute absolute () distance between the prediction and the labels, averaged across the labels. This is normalized by the largest value among the gold labels to normalize it to range.
The overall score is an average of all the above responses. Given that the field distribution of each type does not vary from one model to another, averaging these different measures of quality is acceptable when comparing different models.
3.6 Evaluation Protocol
Our evaluation protocol consistent of a loop over the various choices of evaluation tasks. Specifically means that, the evaluation script provide the model with the URL where the task can be reached. Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
Furthermore, in this paper we also provide the list of field names on which we will evaluate model responses. This arguable simplifies the task a bit, however, our benchmarks leaves it open for any future analysis with open-ended interaction without any additional hand-holding about the names of the target fields.
Task splits for measuring generalization to unseen instructions.
To accommodate for scenarios that may involve fine-tuning on supervised data, we create different task splits: (i) Train: a set of 121 tasks that can be used for supervised model development. (ii) Test: a set of 20 tasks that be used for evaluating model quality. (iii) Test: a set of 17 tasks that are strictly more difficult than the rest of the tasks, because they require more complex combination of actions or other tasks not covered in Table 3. In this version of the work, we do not provide any experiments on Test since they require more sophisticated engineering that is beyond the scope of this work and should be addressed in future work.
4 Evaluating Models in Solving Web-based Tasks in TurkingBench
We benchmark various models with architectural choices on TurkingBench. We note that, our goal here is to provide reasonable baselines for our proposed benchmark and hence, we avoid specialized models that may use specific assumptions or supervision of our data.
Models.
As discussed earlier, our models need to consume information on a mix of information modalities. A model can, therefore, consume the instructions as text, image, or a combination of both. We experiment with state-of-the-art models like GPT-4 and Claude-2.1.666Accessed via their APIs in January 2024. For GPT-4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information. When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page. (2) “relevant” indicates a few lines of HTML above and under the input field the model is currently solving. To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation task. This adds up to web pages evaluated with a total of roughly input fields evaluated.
We note that, besides these general-purpose models there are more specialized text-based models that are specialized for web-exploration either through pre-training on text data (Aghajanyan et al., 2021; 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2022; Chen et al., 2022) or visual perception of web (Dosovitskiy et al., 2020; Rust et al., 2022; Lee et al., 2022). We delegate such explorations that are likely to yield better results to the future.
An oracle model for ceiling performance.
We implement an oracle baseline that mimics the gold labels for each input. This data provides granular action sequences that facilitate future model developments. These action sequences are extracted by connecting crowdworker annotations to the name of the corresponding HTML input elements in task web pages and executing an appropriate action similar to the ones shown in Figure 5. A pseudo-code of the oracle model is shown in Alg.2.777See the a screencast of the oracle model’s execution. We note that, the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) that were included in Test (discussed in our evaluation split §3.5).
A “do-nothing” model as a lower-bound.
We also evaluate a trivial baseline that executes nothing (hence the name, “do-nothing”). As we see later in the results, there are tasks that score more that zero. This is because there are tasks that ask, for example, corrections on a given text where at times doing nothing is the right action. Despite such settings, we will see the overall there is a big margin between the do-nothing baseline and the oracle baseline.
4.1 Empirical Results
We outline our empirical observations on our proposed benchmark.
Modalities |
Score (%) | |||||||
| Do-nothing | – | – | – | – | 7.8 | |||||||
| GPT4 | T | relevant | 7 | 5k | 21.3 | |||||||
| T | full | 7 | 82k | 39.3 | ||||||||
| GPT4-V | T+ V | relevant | 7 | 8.4k | 24.5 | |||||||
| T+ V | full | 7 | 86k | 41.7 | ||||||||
| Claude2.1 | T | full | 7 | 82k | 22.6 | |||||||
| Oracle | – | – | – | – | 100.0 |
Despite the remarkable performance of generalist models, they remain far from our ceiling performance.
We show the results of evaluating our main models in Table 4. For each row, we indicate whether the input to the model contains text-only (T) vs. text-vision (T+V). All the models show have 7 in-context demonstrations of tasks and desired action. Note in the table we also show the size input prompts measured in subwords. The best performance is obtained by GPT-4 vision-language model (T+V) with access to full text of each task. This configuration also happens to have a extremely large prompt length (86k subwords) and it shows the remarkable ability of this model to exploit long-range dependencies. We note that the gains of the vision model (T+V) over text-only model (T) is minimal ( vs. ).
Text-only and vision-language models show complementary capability.
For our top configurations in Table 4 (GPT-4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6. On 12 evaluation tasks (out of 20) the performance of these two models are notably different from one another. While it is intuitive to speculate that task with richer task interface likely benefit more from visual, in practice it was not clear what exact details details about task design leads to such differences.
The gains of in-context demonstrations plateau quickly.
As mentioned, we use few-shot prompting of models in order to steer them their predictions. Here we study the effect of the number of demonstrations in model performance. As the results are shown in Figure 7, the gains of in-context demonstrations quickly plateaus when the number of demonstrations just above 3 demonstrations.
Performance across field type.
To better understand the difficulty of the task as a function of field types, we show the performance of GPT-4 (text input; 7 demonstrations) per input field type in Figure 8. We distinguish between “text” and “text area” because they’re defined with different tags in HTML (<input type="text"> vs. <textarea>). The former is often for short inputs and the latter for longer, multi-sentence outputs. This is also refelcted in lower performance results for “textarea” inputs.
5 Conclusion
We introduced TurkingBench to facilitate research on general-purpose web-based agents. Our benchmark is focused on tasks defined in the context of web pages, such as those that tend to appear on crowdsourcing platforms. Our benchmark includes a comprehensive Python-based framework that facilitates evaluation and model development. We hope this benchmark facilitates further development of web-based assistant agents.
Limitations.
There are various limitations in our work. (i) Despite the complexity of crowdsourcing tasks, they make up just a tiny fraction of the web. We are willing to make this trade-off to obtain natural tasks with ground-truth labels. Other existing benchmarks often need to develop artificial (manually-annotated) tasks and manually annotate gold labels for them. (ii) Because of the choice of domain, our tasks do not involve interactions that require jumping over multiple web pages—instead, all the tasks are contained within a given web page. Despite these simplifications, as we see in our experiments, the task remains challenging for the models. (iii) Our evaluation setup provides some hand-holding for the models by providing the names of the fields that need to be modified. Future work should study variants of our task where such hints are not provided to the model.
Author Contributions
KX led the major implementation of the evaluation framework and obtaining the final results. YK implemented the first version of the repository. KS, YW, AB participated in meetings, provided feedback and helped with the draft. JZ helped with the paper draft. DK lead the project, implementation and writing.
Acknowledgements
This project is supported by a generous gift from Amazon. The authors would like to thank Chris Callison-Burch, Mark Dredze, Karthik Narasimhan, Nikhil Sharma, Owen Bianchi, and Elizabeth Salesky for their constructive discussions. GPU machines for conducting experiments are provided by the ARCH (Rockfish) cluster (https://www.arch.jhu.edu).
6 Ethics Statement
A notable portion of humans spend more time on the web than the outside world (Perrin & Kumar, 2019). Consider an academic individual who spend long hours of their day responding to emails, writing proposals, scheduling talks, setting up meetings, etc. As a result, AI-driven automation of web navigation has the potential to simplify the experience of many and make web more accessible (Deng, 2023).
We acknowledge a concern that this work may enable technologies that would replace crowd workers, who have been an essential part of the AI development lifecycle. The concern is not specific to this work but applies broadly to all AI developments. We note that all the models discussed here are mainly enabled by developments outside this benchmark. Furthermore, the experimental results on our benchmark show that state-of-the-art self-supervised models are still far from the benchmark’s ceiling performance, even on our (simplified) evaluation setup. This means it will take much more progress before AI systems can fully solve autonomous crowdsourcing.
We hope this work enables benevolent use-cases of AI, such as AI systems that help improve the quality and productivity of crowd workers.

### v3

@@
-Tur[k]ingBench: A Challenge Benchmark for Web Agents
+Instructions for *ACL Proceedings Learning to Tackle and Generalize to TurkingBench Can Self-Supervised Models Learn to Tackle TurkingBench?
 Abstract
-Recent chatbots have demonstrated impressive ability to understand and communicate in raw-text form.
-However, there is more to the world than raw text.
-For example, humans spend long hours of their time on web pages, where text is intertwined with other modalities and tasks are accomplished in the form of various complex interactions.
-Can state-of-the-art multi-modal models generalize to such complex domains?
-To address this question, we introduce TurkingBench, a benchmark of tasks formulated as web pages containing textual instructions with multi-modal context.
-Unlike existing work which employs artificially synthesized web pages, here we use natural HTML pages that were originally designed for crowdsourcing workers for various annotation purposes.
-The HTML instructions of each task are also instantiated with various values (obtained from the crowdsourcing tasks) to form new instances of the task.
-This benchmark contains 32.2K instances distributed across 158 tasks.
-Additionally, to facilitate the evaluation on TurkingBench, we develop an evaluation framework that connects the responses of chatbots to modifications on web pages (modifying a text box, checking a radio, etc.).
-We evaluate the performance of state-of-the-art models, including language-only, vision-only, and layout-only models, and their combinations, on this benchmark.
-Our findings reveal that these models perform significantly better than random chance, yet considerable room exists for improvement.
-We hope this benchmark will help facilitate the evaluation and development of web-based agents.111Dataset and baselines are available at: https://turkingbench.github.io
+Can advanced multi-modal models effectively tackle complex web-based tasks?
+Such tasks are often found on crowdsourcing platforms, where crowdworkers engage in challenging micro-tasks within web-based environments.
+Building on this idea, we present TurkingBench, a benchmark consisting of tasks presented as web pages with textual instructions and multi-modal contexts.
+Unlike previous approaches that rely on artificially synthesized web pages, our benchmark uses natural HTML pages originally designed for crowdsourcing workers to perform various annotation tasks.
+Each task’s HTML instructions are instantiated with different values derived from crowdsourcing tasks, creating diverse instances.
+This benchmark includes 32.2K instances spread across 158 tasks.
+To support the evaluation of TurkingBench, we have developed a framework that links chatbot responses to actions on web pages (e.g., modifying a text box, selecting a radio button).
+We assess the performance of cutting-edge private and open-source models, including language-only and vision-language models (such as GPT4 and InternVL), on this benchmark.
+Our results show that while these models outperform random chance, there is still significant room for improvement.
+We hope that this benchmark will drive progress in the evaluation and development of web-based agents.111Dataset and code: https://turkingbench.github.io
 1 Introduction
-There has been tremendous progress in building diverse AI models, thanks to the availability of large pre-trained language models (LLM) (Radford et al., 2019).
-However, the bulk of the progress has been limited to chatbots that process raw text-form in their input or output.
-However, humans rarely work with text-only interfaces.
-Instead, they generally like to work with visually situated language, such as web-based applications where language is intertwined with other visual elements such as images and tables.
-It is only recently that multi-modal chatbots have entered the arena (OpenAI, 2023; Anil et al., 2023; Kil et al., 2024), and plenty is yet to be explored in terms of their capabilities and limits, particularly in the context of web-based exploration.
-A key limitation is the lack of well-defined and natural benchmarks that encompass the complex, interactive, and multi-modal reasoning for various web-based tasks that provide an objective measure of progress for web-navigating agents.
-In this work, we introduce TurkingBench, a benchmark of 158 web-grounded tasks.
-Examples of these tasks are shown in Figure 1.
-The information content in these tasks is relayed via various stylistic elements (colors, font sizes, shapes, etc.), structural knowledge (tables, paragraph headings, etc.), as well as modalities (text, images, videos, etc.) to communicate various complex tasks for human users.
-A thorough understanding of these multi-modal signals is a complex task that requires various skills, including understanding language situated in rich visual cues.
-The tasks in TurkingBench are collected from tasks published on the Amazon Mechanical Turk (AMT) crowdsourcing platform, a well-established platform for connecting human workers to solve micro-tasks in return for a reward.
-We are inspired by the ability humans to solve a variety of tasks through web-based interfaces (Efrat & Levy, 2020).
-There are also other distinguishing factors to TurkingBench in comparison to the existing literature.
-The instructions for solving our tasks are interleaved within each web page, and solving them require effective understanding of the page content.
-This is distinct from most related benchmark (Deng et al., 2023; Zhou et al., 2023) where the task instructions are provided as as a free-form sentences.
-Regarding the naturalness of the data, the web-pages in TurkingBench were originally tailored for human crowd workers, making them more natural compared to the recent benchmarks that engineer simplified web pages (Liu et al., 2018; Furuta et al., 2023).
-Last but not least, since the tasks were originally intended for solving non-trivial annotation tasks on crowdsoucing platforms (e.g., building complex NLP benchmarks) they inherently posses a high degree of complexity.
-We discuss more related work in §2.
-There are a variety of challenges involved in solving our proposed task: (i) Multi-modality—these tasks require the models to “see” or “hear” the task content.
-(ii) Interaction— The agents’ actions are executed in interaction with the web pages, taking into account the context of the previous changes.
-(iii) Long context—the web pages (the HTML encoding, screenshots, etc.) are typically longer than the context window of existing models.
-We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1).
-We observe that GPT4 can perform notably higher than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page.
-However, GPT4’s performance is notably far from the estimated upper bound of the dataset.
-We also analyze model performance as a function of field type and task length, which identify challenges to our future progress.
-In summary, we introduce TurkingBench, a benchmark of web-based tasks that requires multi-modal and interactive reasoning on web pages.
-We formalize the evaluation framework for interacting with and modifying these pages.
-The introduction of this benchmark aims to encourage the development of agent models that are not only capable of understanding language within a multi-modal context, but also able to transform their knowledge into a sequence of affordances when interacting with web pages.
-We evaluate notable models on this benchmark and show that they have noteworthy room for progress.
-We hope this benchmark motivates and enables further development of assistive, web-based agents.
-| MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2023) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
+Significant progress in AI model development has been fueled by large pre-trained language models (LLMs) (Radford et al., 2019; Dubey et al., 2024).
+However, humans usually engage with visually rich environments, like web-based applications that combine language with visual elements such as images and tables.
+Notably, many people now spend more time online than in the physical world (Perrin and Kumar, 2019).
+Crowdsourcing platforms are a key domain where human workers complete micro-tasks in exchange for rewards.
+Each day, hundreds of requesters publish tasks that need annotation by workers.
+These tasks are typically presented as web pages, incorporating stylistic features and structured information (e.g., tables, colors, font sizes) to communicate the tasks effectively.
+Inspired by humans’ ability to solve various tasks through web interfaces (Efrat and Levy, 2020), and with the growing capabilities of multi-modal chatbots (OpenAI, 2023; Anil et al., 2023) that can generate and interact with external tools Chen et al.
+(2021); Dubey et al.
+(2024); Schick et al.
+(2023), we are motivated to assess their performance on the same web-based crowdsourcing tasks that humans handle daily.
+We present TurkingBench, a benchmark for web-grounded tasks.
+Examples of these tasks can be seen in Figure 1.
+The information in these tasks is conveyed through a mix of stylistic elements (such as colors, font sizes, and shapes), structural features (like tables and paragraph headings), and multiple modalities (including text, images, and videos).
+Effectively interpreting these multi-modal signals is a complex challenge that demands a range of skills, particularly the ability to understand language within rich visual contexts.
+The tasks in TurkingBench are drawn from those published on the Amazon Mechanical Turk (AMT) platform, a well-known site where human workers complete micro-tasks for rewards.
+The benchmark comprises 158 web-grounded tasks (web pages), each averaging 16.8K tokens and containing 15.6 input fields that AI models must label or fill in.
+Each task can be instantiated with different input values (e.g., varying questions, paragraphs, images).
+In total, our benchmark includes 36.2K input fields.
+There are several challenges in solving our proposed task: (i) Multi-modality—tasks require models to “see”, “read” or “head”.
+(ii) Interaction—agents’ actions must consider the context of previous decisions on the web pages.
+(iii) Long context—web pages (HTML, screenshots, etc.) often are quite long.
+To facilitate interaction and evaluation of AI systems on these web-based tasks, we developed a Python-based middleware that connects AI models with web pages and also serves as the evaluation framework.
+This development, led by several students, has taken over a year of dedicated effort to ensure smooth functionality.
+MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2024) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
 Inputs | language commands | language commands | customer service queries | shopping queries | language commands | English commands | instructions embedded in web pages |
-| The input construction | crowdsourced | manual | mined online | crowdsourced | author-written & LM-augmented | author-written | collected from MTurk |
-| Environments | simplified web page | simplified but compositional | real-world websites | shopping web pages | real-world websites | variety of web pages | crowdsourcing websites |
+The input construction | crowdsourced | manual | mined online | crowdsourced | crowdsourced | author-written | collected from MTurk |
+Environments/domain | simplified web page | simplified but compositional | real-world websites | shopping web pages | real-world websites | variety of web pages | crowdsourcing websites |
 Natural inputs? | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ |
 Realistic interface? | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ | ✓ |
-| Dynamic interaction | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
+Interaction w/ pages? | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
 Functional correctness? | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✓ |
+Inter-page navigation? | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
+We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1).
+Among our evaluated models, GPT4 performs significantly better than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page.
+Among the open-source models, InternVL-2 achieves the highest performance.
+In both settings, the top-performing models leverage the combination of text and visual modalities, highlighting the advantages of multi-modal approaches.
+Despite these positive gains, GPT4’s performs well below the estimated upper bound of the dataset.
+We analyze model performance based on field type and task length, identifying challenges for future progress.
+Compared to existing web-based benchmarks, TurkingBench fills a critical gap.
+A key feature of our benchmark is that both the web pages and instructions are naturally sourced from the wild, rather than being artificially constructed.
+The web pages in TurkingBench were originally designed for human crowd workers, making them more authentic than recent benchmarks that use simplified, engineered web pages (Liu et al., 2018; Furuta et al., 2023).
+Additionally, task instructions in TurkingBench are embedded within the web pages, requiring a deep understanding of the content to solve them.
+This contrasts with most related benchmarks (Deng et al., 2024; Zhou et al., 2023), where instructions are provided as standalone sentences.
+Finally, because the tasks were initially created for complex annotation on crowdsourcing platforms (e.g., building NLP benchmarks), they inherently carry a high level of complexity.
+We believe this makes TurkingBench a distinctive platform for evaluating both the comprehension and interaction capabilities of LLMs.
+In summary, we present TurkingBench, a benchmark for web-based tasks requiring multi-modal and interactive reasoning.
+We establish an evaluation framework for interacting with and modifying web pages.
+Our evaluation of notable models reveals significant room for improvement.
+We hope this benchmark motivates and aids the advancement of assistive, web-based agents.
 2 Related Work
-Web-navigation benchmarks.
-There are several benchmarks for quantifying abilities of web-navigation agents (Lù et al., 2024; Koh et al., 2024; He et al., 2024).
-Table Table 1 shows a comparison among a subset of these works.
-These benchmarks provide an easy way to evaluate models in a reproducible manner.
-However, all these benchmarks are limited compared to the real-world web.
-For example, most of the existing datasets operate on websites with limited complexity, such as the ones that are pre-specified, synthesized, or simplified (Liu et al., 2018; Li et al., 2020; Yao et al., 2022; Furuta et al., 2023; Cheng et al., 2023).
-In almost all the existing benchmarks, the task definitions (instructions) are done after collecting the websites, through manual crowdsourcing or written manually.
-However, in TurkingBench, instructions are also naturally mined, as they are already embedded in the crowdsourcing web pages.
-It is also worth noting a limitations of TurkingBench compared to other benchmarks: here the tasks are limited to one page (no page navigation).
-Given these trade-offs, the field is likely to benefit from all the aforementioned benchmarks.
-Instruction-following datasets.
-Being able to execute actions on the web also connects to instruction-following literature.
-Currently, the dominant majority of this literature has has focused on raw text instructions and outputs (Wang et al., 2022; Mishra et al., 2022; Bach et al., 2022; Dwivedi-Yu et al., 2022; Ouyang et al., 2022; Sanh et al., 2022; Shen et al., 2023).
-The problem of web-exploration (and the proposed benchmark here) is a strict extension of this literature with multiple modalities, with large inputs (long HTML or large screenshots) and strategic decision-making.
+Several benchmarks exist for evaluating web-navigation agents (Lù et al., 2024; Koh et al., 2024; He et al., 2024; Liu et al., 2024; Pan et al., 2024; Cheng et al., 2023; Zheng et al., 2024b) (see Table 1).
+These benchmarks offer reproducible model assessments but vary in their construction and coverage of real-world web navigation.
+Most focus on websites with limited complexity, such as pre-specified, synthesized, or simplified web pages (Liu et al., 2018; Li et al., 2020; Yao et al., 2022; Furuta et al., 2023).
+In nearly all existing benchmarks, task definitions are created after collecting the websites, typically through manual crowdsourcing.
+In contrast, TurkingBench features more natural task instructions, originally intended for human users on crowdsourcing web pages.
+While this provides more natural task definitions, TurkingBench is limited to tasks involving data annotation that require effective manipulation of complex web pages.
+Given the complementarity of these benchmarks, the research field is likely to benefit from all.
+Unlike our focus on web pages, it is worth noting the benchmarks that utilize other environments for interactive problem solving.
+For instance, benchmarks that concentrate on mobile operating systems such as Android Li et al.
+(2020); Toyama et al.
+(2021); Sun et al.
+(2022); Burns et al.
+(2022) or computer applications in0 operating system environments Trivedi et al.
+(2024); Xie et al.
+(2024).
 Multi-modal reasoning tasks.
-TurkingBench is also related to the efforts on multi-modal reasoning.
-Some works focus on grounding commands to web page elements, setting interaction aside (Pasupat et al., 2018).
-Several recent efforts study the problem in interactive settings (Gur et al., 2018; Ku et al., 2020; Li et al., 2020; 2022; Li & Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021), though their instructions are simple, typically only a few sentences long.
-Our work differs since our instructions are part of the web pages, and inferring them involves multi-modal reasoning.
+TurkingBench is also related to efforts in multi-modal interactive environments (Gur et al., 2018; Ku et al., 2020; Li et al., 2020, 2022; Li and Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021).
+However, these often feature simple instructions that are only a few sentences long, unlike our more extensive instructions embedded within web pages.
 Web-based agents.
-The idea of intelligent automated assistant agents that collaborates with humans to complete tasks has been around (Allen et al., 2007).
-One can view web-based agents as extensions of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022).
-However, general-purpose web-browsing necessitates dealing with challenges well beyond the classic semantic parsing; for example, the diversity of web pages (no two pages are identical), their change over time, the vast action space (a variety of ways that one can interact with web), multi-modality (text, vision, audio), interactive nature of web-browsing, and so on.
-The existing literature use a variety of supervision protocols for supervision, for example, behavior cloning (supervised learning of actions) (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022), in-context prompting and demonstrations (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023).
-Here are a few highlighted works, among others: (Sodhi et al., 2023) proposed web agents equipped with hierarchical policies (HeaP), i.e., low-level actions that are guided by high-level plans.
-(Liu et al., 2023) benchmarks a variety of open-source and open-source and closed-source (API-based) models on several well-known interactive benchmarks.
-While the focus of our work is introducing a new benchmark, we have intended to provide a diverse variety of models to better understand its difficulty.
-Our efforts are not exhaustive and leave more room for future work to explore more effective model engineering.
-3 TurkingBench: Benchmarking Web Agents with Multi-Modal, Interactive Turking Tasks
-We propose the development of TurkingBench.
-The first purpose of this benchmark is to serve as a standardized evaluation of model generalizability.
-In addition to supporting evaluation, a subset of this data can streamline model design and development.
+The concept of intelligent automated assistant agents collaborating with humans to complete tasks has been around for some time (Allen et al., 2007) and can be seen as an extension of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022).
+Recent literature has explored various forms of supervision, including behavior cloning of actions (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022; Liu et al., 2023b), and in-context learning (ICL) (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023).
+Our work focuses on introducing a new benchmark, providing ICL baselines, and leaving the exploration of more sophisticated models for future research.
+3 TurkingBench: Benchmarking Web Agents via Multi-Modal Turking Tasks
+We discuss the development of TurkingBench.
+The primary purpose of this benchmark is to provide a standardized evaluation for web-based agents.
+Additionally, a subset of this data can facilitate model design and development.
+The overall development has undergone several rounds of meticulous engineering, led by multiple students, and has taken over a year of dedicated effort to ensure smooth functionality.
 3.1 Benchmark Schema
-TurkingBench consists of a collection of tasks where each task is a bundle of the following components:
-A web template containing instructions for the task, input variables, and input fields for the outputs.
-Input values to be instantiated for the input variables.
-Annotated output labels provided by crowd workers.
+TurkingBench consists of a collection of tasks where each task is a bundle of the following components: (A) A web template containing instructions for the task, input variables, and input fields for the outputs.
+(B) Input values to be instantiated for the input variables.
+(C) Annotated output labels provided by crowd workers.
 An example is shown is shown in Figure 2.
@@
 For each field, we have values previously collected from crowd workers that we will use for evaluation.
-3.2 Collecting the tasks
-For a benchmark of tasks grounded in web pages to effectively measure progress and generalizability, it needs to be diverse and with broad coverage.
-Toward such a recourse, we plan on utilizing two existing sources: (A) We will collect the crowdsourcing templates and the resulting annotations from various AI researchers who are active requesters on AMT.
-(B) Another incidental resource is the set of tasks published on AMT sandbox,222https://requestersandbox.mturk.com a testing environment for requesters to test out the prototypes of their crowdsourcing templates.
-Since the sandbox environment contains tasks from all users of AMT, it offers more variety than (A).
-However, they would require additional work to annotate and clean up.
+3.2 Collecting the web-based tasks
+For a benchmark of tasks grounded in web pages to effectively measure progress and generalizability, it needs to be diverse and broadly covered.
+The majority of the tasks in TurkingBench are sourced from prior crowdsourcing tasks conducted by the authors and their collaborators over the years.
+We also considered using tasks from the AMT sandbox,222https://requestersandbox.mturk.com a testing environment for requesters to prototype their crowdsourcing templates.
+While this subset could have added more diversity to our data, it would have required additional annotation and cleanup, which we did not pursue.
+During the selection process, we did not restrict ourselves to any specific task type.
+Our collected web pages encompass a variety of tasks, and this diversity is a strength.
+The wide range of tasks captures an array of problems that a generalist agent would encounter, mirroring real-world usability where tasks vary greatly.
+We conducted light quality control to ensure the validity of the instructions and their annotations.
+Here are the steps we took: (1) Several task instructions appeared underdefined (pilot experiments for crowdsourcing tasks), so we eliminated them.
+(2) Certain tasks were missing data artifacts, such as images or videos.
+We eliminated these tasks if the missing media file was critical for solving the task.
+(3) Some tasks did not render properly.
+A subset used older MTurk design conventions that are no longer supported.
+We manually revised or eliminated these tasks.
+(4) Finally, we manually spot-checked the task annotations to ensure the quality was not noisy.
 Ultimately, we put together a collection of 158 crowdsourcing tasks (examples in Figure 1).
@@
 3.3 Statistics
-Figure 3 shows the overall statistics of TurkingBench.
+Table 2 shows the overall statistics of TurkingBench.
 The dataset contain a large sum of input-output instances (36.2K) across 158 tasks.
@@
 3.4 Programmatic Interaction with Web
-We developed a Python library (Figure 4) to streamline the interaction of self-supervised models with our tasks and their evaluations (Figure 4).
+We developed a Python library to streamline the interaction of self-supervised models with our tasks and their evaluations (Figure 4).
 Our library contains various components.
@@
 Therefore, we build our wrappers around these actions to build an action library with enough complexity to cover the dominant majority of tasks in our benchmark.
-| Actions | Modality | Description |
-| modify_text | T | modifies the text of input box |
-| modify_checkbox | T | modifies the selection of checkbox |
-| modify_radio | T | modifies a radio button |
-| modify_select | T | selects an item in a drop-down menu |
-| modify_range | T | modifies a range input |
-| get_html | T | fetches the HTML content of a page |
-| capture_screen | V | fetches the screenshot of a page |
-| click | V | clicks on a given coordinate |
-| type | V | keyboard type at the selected |
-| scroll | – | scrolls up or down |
-| maximize | – | maximizes the web page |
+Actions (Modality) | Description |
+modify_text (T) | modifies the text of input box |
+modify_checkbox (T) | modifies the selection of checkbox |
+modify_radio (T) | modifies a radio button |
+modify_select (T) | selects an item in a drop-down menu |
+modify_range (T) | modifies a range input |
+get_html (T) | fetches the HTML content of a page |
+capture_screen (V) | fetches the screenshot of a page |
+click (V) | clicks on a given coordinate |
+type (V) | keyboard type at the selected |
+scroll (N/A) | scrolls up or down |
+maximize (N/A) | maximizes the web page |
 The list of the actions included in our library is shown in Table 3.
@@
 (iii) Sensing actions allow models to retrieve the latest information about the web page, for example, by fetching its HTML code, taking screenshots of the active page, or around a given element.
+While our actions are designed to cover a broad set of web-based interactions, we acknowledge these are only a subset of what is needed for general-purpose navigation.
+Actions such as “drag-and-drop” are not included here since we did not have any tasks necessitating such actions.
 Web interaction as “tool” resolution.
-The overall task defined here can be viewed as iterative tool resolution based on the context of each task.
-A web agent solving this task must use a sequence of actions from our action library, based on the context of the task’s instruction (see an example in Figure 5).
-This connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2023; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023) that use Tools to ground LM generations in physical environments.
-The set of actions defined in our action library is the set of “tools” that can be executed to obtain specific outcomes.
-The core of this setup is built upon an LM aware of the tools, the problem context, and the ultimate goal.
+The overall task can be viewed as iterative tool resolution based on the context of each task.
+A web agent solving this task must use a sequence of actions from our action library, guided by the task’s instructions (see an example in Figure 5).
+This approach connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2024; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023; Lu et al., 2024), which use tools to ground LM generations in physical environments.
+The actions in our library serve as the “tools” that can be executed to achieve specific outcomes.
+The core of this setup is an LM aware of the tools, the problem context, and the ultimate goal.
 3.5 Evaluation Metrics
-We devise an evaluation metric that is sensitive to the output type:
-Text fields: For a given text response, we compute ROUGE metric (Lin, 2004) against each label and compute their maximum value.
-Radio/select fields: We compute an exact match of the predicted label against the majority vote of gold labels.
-Checkbox fields: The results of the checkbox selection is a set. Therefore, we quantify this as intersection over union metric between the two sets (set of predicted labels and the gold labels).
-Range fields: The range predictions provide real or integer values.
+We devise an evaluation metric that is sensitive to the output type: (i) Text fields: For a given text response, we compute ROUGE metric (Lin, 2004) against each label and compute their maximum value.
+(ii) Radio/select fields: We compute an exact match of the predicted label against the majority vote of gold labels.
+(iii) Checkbox fields: The results of the checkbox selection is a set.
+Therefore, we quantify this as intersection over union metric between the two sets (set of predicted labels and the gold labels).
+(iv) Range fields: The range predictions provide real or integer values.
 We compute absolute () distance between the prediction and the labels, averaged across the labels.
@@
 Our evaluation protocol consistent of a loop over the various choices of evaluation tasks.
-Specifically means that, the evaluation script provide the model with the URL where the task can be reached.
+Specifically means that, the evaluation script provides the model with the URL where the task can be reached.
 Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
-Furthermore, in this paper we also provide the list of field names on which we will evaluate model responses.
-This arguable simplifies the task a bit, however, our benchmarks leaves it open for any future analysis with open-ended interaction without any additional hand-holding about the names of the target fields.
+Based on this evaluation protocol, it should be evident that our benchmark does not require navigation between web pages, as noted in the limitations section.
+However, our benchmark does require multiple rounds of interaction to solve a given task.
+This is because nearly all of our tasks involve more than one step (input field), each needing to be addressed in a different round of interaction.
+To reduce the difficulty of the task for our models, we provide the list of field names on which we will evaluate model responses.
+While this arguably simplifies the task, our benchmark remains open for future analysis with open-ended interaction, without any additional guidance on the names of the target fields.
 Task splits for measuring generalization to unseen instructions.
-To accommodate for scenarios that may involve fine-tuning on supervised data, we create different task splits: (i) Train: a set of 121 tasks that can be used for supervised model development.
-(ii) Test: a set of 20 tasks that be used for evaluating model quality.
+To accommodate for scenarios that may involve fine-tuning on supervised data, we create different task splits: (i) Train: a set of 125 tasks that can be used for supervised model development.
+(ii) Test: a set of 16 tasks that be used for evaluating model quality (names shown in the caption of Fig.
+6).
 (iii) Test: a set of 17 tasks that are strictly more difficult than the rest of the tasks, because they require more complex combination of actions or other tasks not covered in Table 3.
@@
 4 Evaluating Models in Solving Web-based Tasks in TurkingBench
-We benchmark various models with architectural choices on TurkingBench.
-We note that, our goal here is to provide reasonable baselines for our proposed benchmark and hence, we avoid specialized models that may use specific assumptions or supervision of our data.
+We benchmark various models with different architectures on TurkingBench.
+Our goal is to provide reasonable baselines for our proposed benchmark, so we avoid specialized models that rely on specific assumptions or supervision of our data.
 Models.
@@
 A model can, therefore, consume the instructions as text, image, or a combination of both.
-We experiment with state-of-the-art models like GPT-4 and Claude-2.1.666Accessed via their APIs in January 2024.
-For GPT-4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information.
+We experiment with state-of-the-art proprietary models like GPT4 and Claude-2.1.666GPT4/GPT4-V/Claude accessed via APIs in January 2024; GPT4o accessed in August 2024.
+For GPT4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information.
+We compare the performance of these models with the latest open-source vision language models like LLaVA-1.6 (Liu et al., 2023a), InternVL2 (Chen et al., 2024) and text-only models like Llama-3.1-Instruct Dubey et al.
+(2024).
+Our evaluation focused on major model families (GPT4, LLaVA1.6, etc.), as our goal was setting baselines with widely-recognized models.
+We note that, besides these general-purpose models, there are more specialized text-based models for web exploration, either through pre-training on text data (Aghajanyan et al., 2021, 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2023; Chen et al., 2022), or visual perception of the web (Dosovitskiy et al., 2021; Rust et al., 2023; Lee et al., 2022; Kil et al., 2024).
+We leave such explorations, which are likely to yield better results, to future work.
+Encoding the tasks for evaluation.
 When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page.
-(2) “relevant” indicates a few lines of HTML above and under the input field the model is currently solving.
-To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation task.
+(2) “relevant” indicates a few neighboring lines of HTML adjacent to (above and under) the input field the model is currently solving.
+To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation tasks.
 This adds up to web pages evaluated with a total of roughly input fields evaluated.
-We note that, besides these general-purpose models there are more specialized text-based models that are specialized for web-exploration either through pre-training on text data (Aghajanyan et al., 2021; 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2022; Chen et al., 2022) or visual perception of web (Dosovitskiy et al., 2020; Rust et al., 2022; Lee et al., 2022).
-We delegate such explorations that are likely to yield better results to the future.
-An oracle model for ceiling performance.
+Modalities |
+Score (%) | ||||||||||||
+Do-nothing | – | – | – | – | – | – | 7.8 | ||||||||||||
+open-source models | Llama3.1-Instruct | 8B | T | relevant | 3 | 1k | 128k | 23.2 | |||||||||||
+8B | T | relevant | 7 | 5k | 128k | 25.0 | |||||||||||||
+[…]
+[…]
+13B | T+ V | relevant | 3 | 1.6k | 4096 | 19.8 | |||||||||||||
+[…]
+76B | T+ V | relevant | 3 | 1.6k | 8192 | 30.4 | |||||||||||||
+closed-source models | Claude2.1 | Unknown | T | full | 7 | 82k | 200k | 22.6 | |||||||||||
+[…]
+T | relevant | 3 | 4k | 128k | 18.4 | ||||||||||||||
+T | relevant | 7 | 5k | 128k | 21.3 | ||||||||||||||
+[…]
+T | full | 3 | 28k | 128k | 35.7 | ||||||||||||||
+T | full | 7 | 82k | 128k | 39.3 | ||||||||||||||
+[…]
+T+ V | relevant | 3 | 1.6k | 128k | 30.3 | ||||||||||||||
+T+ V | relevant | 7 | 8k | 128k | 30.19 | ||||||||||||||
+[…]
+T+ V | full | 3 | 30k | 128k | 41.1 | ||||||||||||||
+T+ V | full | 7 | 86k | 128k | 41.7 | ||||||||||||||
+Oracle | – | – | – | – | – | – | 100.0 |
+An oracle for ceiling performance.
 We implement an oracle baseline that mimics the gold labels for each input.
-This data provides granular action sequences that facilitate future model developments.
-These action sequences are extracted by connecting crowdworker annotations to the name of the corresponding HTML input elements in task web pages and executing an appropriate action similar to the ones shown in Figure 5.
-A pseudo-code of the oracle model is shown in Alg.2.777See the a screencast of the oracle model’s execution.
-We note that, the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) that were included in Test (discussed in our evaluation split §3.5).
+This oracle agent ensures the functional correctness of our evaluation.
+Our end-to-end evaluation process (model output execution lookup answers from web pages evaluation) is significantly more complex than a typical NLP benchmark, and any of these steps can fail.
+Achieving 100% accuracy with the oracle (ensuring functional correctness) took the lead student several months of effort.
+The oracle baseline essentially replicates the action sequences of crowdworkers for each HTML input element in the task web pages, executing appropriate actions similar to those shown in Figure 5.
+A pseudo-code of the oracle model is shown in Pseudocode 2.
+We note that the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) included in Test (discussed in our evaluation split §3.5).
+A future use case of this oracle baseline can be to obtain granular action sequences for supervising models.
 A “do-nothing” model as a lower-bound.
-We also evaluate a trivial baseline that executes nothing (hence the name, “do-nothing”).
-As we see later in the results, there are tasks that score more that zero.
-This is because there are tasks that ask, for example, corrections on a given text where at times doing nothing is the right action.
-Despite such settings, we will see the overall there is a big margin between the do-nothing baseline and the oracle baseline.
+We evaluate a trivial baseline that performs no action (no actions (hence, “do-nothing”).
+As we see in the results, this baseline scores more than zero because on some tasks doing nothing is the right action (e.g., making grammatical corrections to a given text that happens to be grammatical).
 4.1 Empirical Results
-We outline our empirical observations on our proposed benchmark.
-Modalities |
-Score (%) | |||||||
-| Do-nothing | – | – | – | – | 7.8 | |||||||
-| GPT4 | T | relevant | 7 | 5k | 21.3 | |||||||
-| T | full | 7 | 82k | 39.3 | ||||||||
-| GPT4-V | T+ V | relevant | 7 | 8.4k | 24.5 | |||||||
-| T+ V | full | 7 | 86k | 41.7 | ||||||||
-| Claude2.1 | T | full | 7 | 82k | 22.6 | |||||||
-| Oracle | – | – | – | – | 100.0 |
+We present the results of evaluating our main models in section 4.
+We experimented with models of varying parameter sizes.
+For each row, we indicate whether the model input contains text-only (T) or text-vision (T+V).
+Additionally, the table shows the size of input prompts measured in GPT-2 subwords Radford et al.
+(2019).
+Wherever possible, we evaluate the models with 7 in-context demonstrations of tasks and desired actions.
+The open-source vision-language models (T+V) had much shorter context windows, so we evaluated them using the “relevant” portion of the HTML code for the task or in-context demonstrations.
 Despite the remarkable performance of generalist models, they remain far from our ceiling performance.
-We show the results of evaluating our main models in Table 4.
-For each row, we indicate whether the input to the model contains text-only (T) vs.
-text-vision (T+V).
-All the models show have 7 in-context demonstrations of tasks and desired action.
-Note in the table we also show the size input prompts measured in subwords.
-The best performance is obtained by GPT-4 vision-language model (T+V) with access to full text of each task.
+The best performance is obtained by GPT4 vision-language model (T+V) with access to full text of each task.
+This is in-line with other recent observations Zheng et al.
+(2024a).
 This configuration also happens to have a extremely large prompt length (86k subwords) and it shows the remarkable ability of this model to exploit long-range dependencies.
@@
 ).
+Open-source models rivaling GPT4.
+Llama3.1-Instruct (8B params) notably outperforms GPT4 (text-only), achieving a score of 25.0% compared to GPT4’s 21.3% with 7 demonstrations when “relevant” bits of the input HTML are supplied.
+This result is particularly impressive given that Llama3.1-Instruct operates with significantly fewer parameters (8B) than GPT4’s rumored size.
+The comparison highlights Llama3.1-Instruct’s ability to efficiently leverage the provided context, potentially making it better suited for certain tasks despite its smaller size.
+Additionally, InternVL2 demonstrates strong performance, coming very close to GPT4 when evaluated with the “relevant” HTML.
+InternVL2’s score of 31.0% is remarkably close to GPT4’s 33%, despite the substantial difference in parameter sizes.
+This near-equivalence underscores the potential of InternVL2 to compete with proprietary models, offering similar performance with a more compact architecture.
+Other open-source models, such as LLaVA-VL-Vicuna/Mistral (7B and 13B parameters), show varying degrees of success, with performance scores ranging from 13.8% to 22.7%.
+While these models are still behind Llama3.1 and InternVL2, their results demonstrate the diverse capabilities and potential of open-source approaches.
 Text-only and vision-language models show complementary capability.
-For our top configurations in Table 4 (GPT-4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6.
-On 12 evaluation tasks (out of 20) the performance of these two models are notably different from one another.
-While it is intuitive to speculate that task with richer task interface likely benefit more from visual, in practice it was not clear what exact details details about task design leads to such differences.
+For our top configurations in section 4 (GPT4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6.
+On 9 evaluation tasks (out of 16) the performance of these two models are notably different from one another.
+While it is intuitive to speculate that task with richer task interface likely benefit more from visual, empirical this it was not obvious what exact details details about task design leads to such differences.
 The gains of in-context demonstrations plateau quickly.
@@
 As the results are shown in Figure 7, the gains of in-context demonstrations quickly plateaus when the number of demonstrations just above 3 demonstrations.
-Performance across field type.
-To better understand the difficulty of the task as a function of field types, we show the performance of GPT-4 (text input; 7 demonstrations) per input field type in Figure 8.
-We distinguish between “text” and “text area” because they’re defined with different tags in HTML (<input type="text"> vs.
+Performance across field types.
+To better understand the task difficulty based on field types, we show the performance of GPT4 (text input; 7 demonstrations) for each input field type in Figure 8.
+We distinguish between "text" and "textarea" fields because they are defined with different HTML tags (<input type="text"> vs.
 <textarea>).
-The former is often for short inputs and the latter for longer, multi-sentence outputs.
-This is also refelcted in lower performance results for “textarea” inputs.
-5 Conclusion
-We introduced TurkingBench to facilitate research on general-purpose web-based agents.
-Our benchmark is focused on tasks defined in the context of web pages, such as those that tend to appear on crowdsourcing platforms.
-Our benchmark includes a comprehensive Python-based framework that facilitates evaluation and model development.
-We hope this benchmark facilitates further development of web-based assistant agents.
+The former is typically used for short inputs, while the latter is for longer, multi-sentence outputs.
+A notable finding was the surprisingly lower performance of GPT4 compared to InternVL2 on “text” input fields, which we could not explain.
+However, since the data is skewed toward checkboxes (Figure 3), GPT4 performs better in the aggregate evaluations.
+4.2 Error Analysis
+We conducted human annotations to better understand the results in section 4.
+This analysis uses predictions from GPT4o (7 demos and “full” encoding) as it is one of the highest-performing models.
+One of the authors reviewed one instance from each of the 16 tasks, totaling 144 responses to the input fields.
+Evaluating GPT4 responses.
+Our annotator directly evaluated 134 (out of 144) responses that we successfully parsed by our evaluation metric (i.e., no parsing error).
+The annotator agreed with GPT4o’s responses in 60% (80 out of 134) of the cases, reaffirming that our benchmark remains challenging for the models.
+Our analysis revealed several recurring issues.
+One common problem occurred in binary classification tasks, such as determining whether a word is an adjective with a similar meaning to a given word.
+For example, GPT4o incorrectly classified “discernible” as describing “modernity,” despite the lack of synonymy.
+Another significant discrepancy was observed in tasks requiring GPT4o to generate both correct and incorrect answers.
+Instead of providing actual incorrect answers, GPT4o sometimes returned placeholders like //incorrect options//, failing to meet the task’s requirements.
+In some cases, GPT4o also made syntax errors.
+However, the evaluation focused on the content of the responses rather than their syntactical correctness, so syntax errors did not affect the assessment of answer accuracy.
+We would like to provide more details upon having more space.
+5 Conclusion and Future Work
+We introduced TurkingBench to facilitate research on web-based agents.
+Our benchmark focuses on tasks defined within the context of web pages, such as those commonly found on crowdsourcing platforms.
+It includes a comprehensive Python-based framework that supports both evaluation and model development.
+We hope this benchmark will drive further advancements in the development of web-based assistant agents.
+Future work should explore modeling improvements for better web agents.
+For instance, a RAG-style approach could semantically chunk web pages into meaningful segments, a non-trivial task.
+Another avenue could involve using these agents as CoPilots for human annotation on crowdsourcing platforms, helping workers identify potential mistakes.
+We consider these extensions somewhat orthogonal to the primary focus of this work and hope future research will address them.
 Limitations.
-There are various limitations in our work.
-(i) Despite the complexity of crowdsourcing tasks, they make up just a tiny fraction of the web.
-We are willing to make this trade-off to obtain natural tasks with ground-truth labels.
-Other existing benchmarks often need to develop artificial (manually-annotated) tasks and manually annotate gold labels for them.
-(ii) Because of the choice of domain, our tasks do not involve interactions that require jumping over multiple web pages—instead, all the tasks are contained within a given web page.
-Despite these simplifications, as we see in our experiments, the task remains challenging for the models.
-(iii) Our evaluation setup provides some hand-holding for the models by providing the names of the fields that need to be modified.
-Future work should study variants of our task where such hints are not provided to the model.
-Author Contributions
-KX led the major implementation of the evaluation framework and obtaining the final results.
-YK implemented the first version of the repository.
-KS, YW, AB participated in meetings, provided feedback and helped with the draft.
-JZ helped with the paper draft.
-DK lead the project, implementation and writing.
+We discuss several limitations: (i) No navigation between pages: Our benchmark does not require navigation between web pages.
+While inter-page navigation is important, it is not the only challenge.
+Effective understanding and manipulation of each page, which is our focus, remains a significant challenge for web agents.
+Our benchmark still requires navigation within each page, involving multiple rounds of interaction to solve tasks with multiple input fields.
+As shown in our experiments, this remains challenging for the models.
+We accept this trade-off to obtain more natural tasks compared to most existing benchmarks.
+(ii) Simplified evaluation: This is not an inherent limitation of our benchmark but a simplifying assumption in our evaluation.
+For simplicity, our evaluation setup provides some guidance by specifying the names of the fields to be modified.
+Future work should explore variants of our experiments where such hints are not provided to the model.
+Ethical considerations.
+We recognize concerns that this work could lead to technologies replacing crowd workers, who are vital to AI development.
+This concern is not unique to our work extends to all AI.
+Our results show that state-of-the-art models are still far from fully automating crowdsourcing tasks, even with simplified evaluation.
+Therefore, we hope our work enables benevolent use-cases of AI such as enhancing the quality and productivity of crowd workers rather than replacing them.
 Acknowledgements
-This project is supported by a generous gift from Amazon.
+This project is partly supported by ONR grant (N00014-24-1-2089) and a generous gift from Amazon.
 The authors would like to thank Chris Callison-Burch, Mark Dredze, Karthik Narasimhan, Nikhil Sharma, Owen Bianchi, and Elizabeth Salesky for their constructive discussions.
 GPU machines for conducting experiments are provided by the ARCH (Rockfish) cluster (https://www.arch.jhu.edu).
-6 Ethics Statement
-A notable portion of humans spend more time on the web than the outside world (Perrin & Kumar, 2019).
-Consider an academic individual who spend long hours of their day responding to emails, writing proposals, scheduling talks, setting up meetings, etc.
-As a result, AI-driven automation of web navigation has the potential to simplify the experience of many and make web more accessible (Deng, 2023).
-We acknowledge a concern that this work may enable technologies that would replace crowd workers, who have been an essential part of the AI development lifecycle.
-The concern is not specific to this work but applies broadly to all AI developments.
-We note that all the models discussed here are mainly enabled by developments outside this benchmark.
-Furthermore, the experimental results on our benchmark show that state-of-the-art self-supervised models are still far from the benchmark’s ceiling performance, even on our (simplified) evaluation setup.
-This means it will take much more progress before AI systems can fully solve autonomous crowdsourcing.
-We hope this work enables benevolent use-cases of AI, such as AI systems that help improve the quality and productivity of crowd workers.

### v4

@@
-Instructions for *ACL Proceedings Learning to Tackle and Generalize to TurkingBench Can Self-Supervised Models Learn to Tackle TurkingBench?
+Tur[k]ingBench: A Challenge Benchmark for Web Agents
 Abstract
@@
 This benchmark includes 32.2K instances spread across 158 tasks.
-To support the evaluation of TurkingBench, we have developed a framework that links chatbot responses to actions on web pages (e.g., modifying a text box, selecting a radio button).
+To support the evaluation of TurkingBench, we have developed a framework that links chatbot responses to actions on web pages (e.g., modifying a text box, or selecting a radio button).
 We assess the performance of cutting-edge private and open-source models, including language-only and vision-language models (such as GPT4 and InternVL), on this benchmark.
 Our results show that while these models outperform random chance, there is still significant room for improvement.
-We hope that this benchmark will drive progress in the evaluation and development of web-based agents.111Dataset and code: https://turkingbench.github.io
+We hope that this benchmark will drive progress in the evaluation and development of web agents.111 Data and code: https://turkingbench.github.io/
+NAACL ’25
+Tur[k]ingBench: A Challenge Benchmark for Web Agents
+Kevin Xu∗ Yeganeh Kordi Tanay Nayak Adi Asija Yizhong Wang Kate Sanders Adam Byerly Jingyu Zhang Benjamin Van Durme Daniel Khashabi††thanks: Correspondence to: {kxu39,danielk}@jhu.edu Johns Hopkins University Brown University University of Washington
 1 Introduction
@@
 We present TurkingBench, a benchmark for web-grounded tasks.
-Examples of these tasks can be seen in Figure 1.
+An example task can be seen in Figure 1.
 The information in these tasks is conveyed through a mix of stylistic elements (such as colors, font sizes, and shapes), structural features (like tables and paragraph headings), and multiple modalities (including text, images, and videos).
@@
 In total, our benchmark includes 36.2K input fields.
-There are several challenges in solving our proposed task: (i) Multi-modality—tasks require models to “see”, “read” or “head”.
+There are several challenges in solving our proposed task: (i) Multi-modality—tasks require models to “see” or “read”.
 (ii) Interaction—agents’ actions must consider the context of previous decisions on the web pages.
@@
 This development, led by several students, has taken over a year of dedicated effort to ensure smooth functionality.
+We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1).
+Among our evaluated models, GPT4 performs significantly better than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page.
+Among the open-source models, Qwen2 achieves the highest performance.
+In both settings, the top-performing models leverage the combination of text and visual modalities, highlighting the advantages of multi-modal approaches.
+Despite these positive gains, GPT4’s performs well below the estimated upper bound of the dataset.
+We analyze model performance based on field type and task length, identifying challenges for future progress.
+We chose crowdsourcing tasks for their diverse web interactions, including multimodal reasoning, complex form filling, and sequential dependencies—common in e-commerce, education, and administration.
+While rooted in this domain, our benchmark’s challenges (e.g., diverse inputs, long-context reasoning, and multimodal interactions) apply broadly.
+Effective page understanding and manipulation remain key challenges for web agents.
+Compared to existing web-based benchmarks, TurkingBench fills a critical gap.
+A key feature of our benchmark is that both the web pages and instructions are naturally sourced from the wild, rather than being artificially constructed.
+The web pages in TurkingBench were originally designed for human crowd workers, making them more authentic than recent benchmarks that use simplified, engineered web pages (Liu et al., 2018; Furuta et al., 2023).
+Additionally, task instructions in TurkingBench are embedded within the web pages, requiring a deep understanding of the content to solve them.
+This contrasts with most related benchmarks (Deng et al., 2024; Zhou et al., 2023), where instructions are provided as standalone sentences.
+Finally, because the tasks were initially created for complex annotation on crowdsourcing platforms (e.g., building NLP benchmarks), they inherently carry a high level of complexity.
+We believe this makes TurkingBench a distinctive platform for evaluating both the comprehension and interaction capabilities of LLMs.
 MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2024) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
@@
 Inter-page navigation? | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
-We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1).
-Among our evaluated models, GPT4 performs significantly better than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page.
-Among the open-source models, InternVL-2 achieves the highest performance.
-In both settings, the top-performing models leverage the combination of text and visual modalities, highlighting the advantages of multi-modal approaches.
-Despite these positive gains, GPT4’s performs well below the estimated upper bound of the dataset.
-We analyze model performance based on field type and task length, identifying challenges for future progress.
-Compared to existing web-based benchmarks, TurkingBench fills a critical gap.
-A key feature of our benchmark is that both the web pages and instructions are naturally sourced from the wild, rather than being artificially constructed.
-The web pages in TurkingBench were originally designed for human crowd workers, making them more authentic than recent benchmarks that use simplified, engineered web pages (Liu et al., 2018; Furuta et al., 2023).
-Additionally, task instructions in TurkingBench are embedded within the web pages, requiring a deep understanding of the content to solve them.
-This contrasts with most related benchmarks (Deng et al., 2024; Zhou et al., 2023), where instructions are provided as standalone sentences.
-Finally, because the tasks were initially created for complex annotation on crowdsourcing platforms (e.g., building NLP benchmarks), they inherently carry a high level of complexity.
-We believe this makes TurkingBench a distinctive platform for evaluating both the comprehension and interaction capabilities of LLMs.
-In summary, we present TurkingBench, a benchmark for web-based tasks requiring multi-modal and interactive reasoning.
-We establish an evaluation framework for interacting with and modifying web pages.
-Our evaluation of notable models reveals significant room for improvement.
-We hope this benchmark motivates and aids the advancement of assistive, web-based agents.
+In summary, we introduce TurkingBench, a benchmark for multimodal, interactive web-based tasks, along with an evaluation framework for web interaction.
+Our analysis of leading models highlights significant gaps, aiming to drive progress in assistive web agents.
 2 Related Work
@@
 (2022); Burns et al.
-(2022) or computer applications in0 operating system environments Trivedi et al.
+(2022) or computer applications in operating system environments Trivedi et al.
 (2024); Xie et al.
 (2024).
-Multi-modal reasoning tasks.
-TurkingBench is also related to efforts in multi-modal interactive environments (Gur et al., 2018; Ku et al., 2020; Li et al., 2020, 2022; Li and Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021).
-However, these often feature simple instructions that are only a few sentences long, unlike our more extensive instructions embedded within web pages.
-Web-based agents.
-The concept of intelligent automated assistant agents collaborating with humans to complete tasks has been around for some time (Allen et al., 2007) and can be seen as an extension of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022).
-Recent literature has explored various forms of supervision, including behavior cloning of actions (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022; Liu et al., 2023b), and in-context learning (ICL) (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023).
-Our work focuses on introducing a new benchmark, providing ICL baselines, and leaving the exploration of more sophisticated models for future research.
 3 TurkingBench: Benchmarking Web Agents via Multi-Modal Turking Tasks
@@
 (C) Annotated output labels provided by crowd workers.
-An example is shown is shown in Figure 2.
+An example is shown in Figure 2.
 As the example shows, the HTML template contains variables (e.g., ${sys10}) that are instantiated with input values.
@@
 These include but are not limited to language processing tasks (such as paraphrasing, validating factuality, entailment, sentiments, classification, dialogue quality, and rationale generation) and tasks that involve processing images or videos.
-3.3 Statistics
+Statistics.
+The dataset contain a large sum of input-output instances (36.2K) across 158 tasks.
 Table 2 shows the overall statistics of TurkingBench.
-The dataset contain a large sum of input-output instances (36.2K) across 158 tasks.
 Furthermore, Figure 3 shows the distribution of the input fields represented in our benchmark.
-The most common field is “radio,” which is expected as it is commonly used to identify preferences.
+The most common field is “checkbox,” which is expected as it is commonly used when multiple options could be true.
 Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
@@
 Web interaction as “tool” resolution.
-The overall task can be viewed as iterative tool resolution based on the context of each task.
-A web agent solving this task must use a sequence of actions from our action library, guided by the task’s instructions (see an example in Figure 5).
-This approach connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2024; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023; Lu et al., 2024), which use tools to ground LM generations in physical environments.
-The actions in our library serve as the “tools” that can be executed to achieve specific outcomes.
-The core of this setup is an LM aware of the tools, the problem context, and the ultimate goal.
+The task here can be framed as iterative tool resolution based on task context.
+A web agent must execute a sequence of executable actions from our library, guided by task instructions (see Figure 1).
+This aligns with tool-augmented LMs (Schick et al., 2023; Lu et al., 2024; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023; Lu et al., 2024), where tools ground LM generations in an environment.
+Here, our library’s actions function as “tools” for achieving specific outcomes, with an LM aware of the tools, task context, and final goal.
 3.4 Evaluation Metrics
@@
 3.5 Evaluation Protocol
-Our evaluation protocol consistent of a loop over the various choices of evaluation tasks.
-Specifically means that, the evaluation script provides the model with the URL where the task can be reached.
-Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
+Our evaluation iterates through different choices of evaluation tasks (Pseudocode 1).
+Specifically, the evaluation script supplies the model with the URL where the task is accessible.
+Additionally, the model has access to our action library, enabling executable actions.
 Based on this evaluation protocol, it should be evident that our benchmark does not require navigation between web pages, as noted in the limitations section.
@@
 (ii) Test: a set of 16 tasks that be used for evaluating model quality (names shown in the caption of Fig.
-6).
+5).
 (iii) Test: a set of 17 tasks that are strictly more difficult than the rest of the tasks, because they require more complex combination of actions or other tasks not covered in Table 3.
 In this version of the work, we do not provide any experiments on Test since they require more sophisticated engineering that is beyond the scope of this work and should be addressed in future work.
+Screen Size Considerations.
+When benchmarking multimodal models, the capture_screen action retrieves visual information from the screen.
+A potential concern is the impact of screen size on benchmarking consistency.
+To address this, all evaluations using capture_screen run in headless mode, where the browser operates without a monitor and uses predefined, consistent virtual screen dimensions.
+This eliminates hardware-specific influences, ensuring robust and reproducible results unaffected by screen size variability.
 4 Evaluating Models in Solving Web-based Tasks in TurkingBench
@@
 Our goal is to provide reasonable baselines for our proposed benchmark, so we avoid specialized models that rely on specific assumptions or supervision of our data.
-Models.
-As discussed earlier, our models need to consume information on a mix of information modalities.
-A model can, therefore, consume the instructions as text, image, or a combination of both.
-We experiment with state-of-the-art proprietary models like GPT4 and Claude-2.1.666GPT4/GPT4-V/Claude accessed via APIs in January 2024; GPT4o accessed in August 2024.
-For GPT4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information.
-We compare the performance of these models with the latest open-source vision language models like LLaVA-1.6 (Liu et al., 2023a), InternVL2 (Chen et al., 2024) and text-only models like Llama-3.1-Instruct Dubey et al.
-(2024).
-Our evaluation focused on major model families (GPT4, LLaVA1.6, etc.), as our goal was setting baselines with widely-recognized models.
-We note that, besides these general-purpose models, there are more specialized text-based models for web exploration, either through pre-training on text data (Aghajanyan et al., 2021, 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2023; Chen et al., 2022), or visual perception of the web (Dosovitskiy et al., 2021; Rust et al., 2023; Lee et al., 2022; Kil et al., 2024).
-We leave such explorations, which are likely to yield better results, to future work.
-Encoding the tasks for evaluation.
-When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page.
-(2) “relevant” indicates a few neighboring lines of HTML adjacent to (above and under) the input field the model is currently solving.
-To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation tasks.
-This adds up to web pages evaluated with a total of roughly input fields evaluated.
 Modalities |
@@
 […]
+3B | T | relevant | 7 | 5k | 128k | 20.3 | |||||||||||||
+[…]
+11B | T+ V | relevant | 7 | 6k | 128k | 17.8 | |||||||||||||
 […]
@@
 […]
-| 76B | T+ V | relevant | 3 | 1.6k | 8192 | 30.4 | |||||||||||||
+40B | T+ V | relevant | 7 | 8k | 8192 | 26.1 | |||||||||||||
+[…]
+76B | T+ V | relevant | 7 | 8k | 8192 | 29.6 | |||||||||||||
+[…]
 closed-source models | Claude2.1 | Unknown | T | full | 7 | 82k | 200k | 22.6 | |||||||||||
@@
 Oracle | – | – | – | – | – | – | 100.0 |
+Models.
+As discussed earlier, our models need to consume information on a mix of information modalities.
+A model can, therefore, consume the instructions as text, image, or a combination of both.
+We experiment with state-of-the-art proprietary models like GPT4 and Claude-2.1.666GPT4/GPT4-V/Claude accessed via APIs in January 2024; GPT4o accessed in August 2024.
+For GPT4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information.
+We compare the performance of these models with the latest open-source vision language models like LLaVA-1.6 (Liu et al., 2023a), InternVL2 (Chen et al., 2024) and text-only models like Llama-3.1-Instruct Dubey et al.
+(2024) and Qwen2 Yang et al.
+(2024).
+Our evaluation focused on major model families (GPT4, LLaVA1.6, etc.), as our goal was setting baselines with widely-recognized models.
+We note that, besides these general-purpose models, there are more specialized text-based models for web exploration, either through pre-training on text data (Aghajanyan et al., 2021, 2022; Gur et al., 2022), specialized models for analyzing web-based components (Huang et al., 2022; Tao et al., 2022; Ebrahimi et al., 2023; Chen et al., 2022), or visual perception of the web (Dosovitskiy et al., 2021; Rust et al., 2023; Lee et al., 2022; Kil et al., 2024).
+We leave such explorations, which are likely to yield better results, to future work.
+Encoding the tasks for evaluation.
+When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page.
+(2) “relevant” indicates a few neighboring lines of HTML adjacent to (above and under) the input field the model is currently solving (further details in §C.) To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation tasks.
+This adds up to web pages evaluated with a total of roughly input fields evaluated.
 An oracle for ceiling performance.
-We implement an oracle baseline that mimics the gold labels for each input.
+We implement an oracle baseline that mimics the gold labels for each input (see the logic in 2).
+777See the a screencast of the oracle’s execution.
 This oracle agent ensures the functional correctness of our evaluation.
@@
 The oracle baseline essentially replicates the action sequences of crowdworkers for each HTML input element in the task web pages, executing appropriate actions similar to those shown in Figure 1.
-A pseudo-code of the oracle model is shown in Pseudocode 2.
 We note that the oracle baseline is limited to tasks that do not require complex annotations (such as drag-and-drop) included in Test (discussed in our evaluation split §3.4).
@@
 The comparison highlights Llama3.1-Instruct’s ability to efficiently leverage the provided context, potentially making it better suited for certain tasks despite its smaller size.
-Additionally, InternVL2 demonstrates strong performance, coming very close to GPT4 when evaluated with the “relevant” HTML.
-InternVL2’s score of 31.0% is remarkably close to GPT4’s 33%, despite the substantial difference in parameter sizes.
-This near-equivalence underscores the potential of InternVL2 to compete with proprietary models, offering similar performance with a more compact architecture.
-Other open-source models, such as LLaVA-VL-Vicuna/Mistral (7B and 13B parameters), show varying degrees of success, with performance scores ranging from 13.8% to 22.7%.
-While these models are still behind Llama3.1 and InternVL2, their results demonstrate the diverse capabilities and potential of open-source approaches.
-Text-only and vision-language models show complementary capability.
-For our top configurations in section 4 (GPT4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6.
-On 9 evaluation tasks (out of 16) the performance of these two models are notably different from one another.
-While it is intuitive to speculate that task with richer task interface likely benefit more from visual, empirical this it was not obvious what exact details details about task design leads to such differences.
-The gains of in-context demonstrations plateau quickly.
-As mentioned, we use few-shot prompting of models in order to steer them their predictions.
+Additionally, Qwen2 (72B) demonstrates strong performance, coming very close to GPT4 when evaluated with the “relevant” HTML.
+Qwen2’s score of 34.1% is not too far off from GPT4-V’s 41.7%, which underscores the potential and promise of open-source models over proprietary models.
+4.2 Analysis: Test vs Vision Models
+For one of our top configurations in section 4 (GPT-4, 7 demonstrations with “full” HTML encoding), we present a breakdown of model performance across tasks in Figure 5.
+On 9 out of 16 evaluation tasks, the two models exhibit notable performance differences.
+While it’s intuitive to assume tasks with richer interfaces benefit more from visual input, empirical results do not clearly pinpoint which task design aspects drive these differences.
+Overall, we conclude that text-only and vision-language models have complementary capabilities.
+4.3 Analysis: Effect of the Number of Demos
+As mentioned earlier, we use few-shot prompting of models in order to steer them their predictions.
 Here we study the effect of the number of demonstrations in model performance.
 As the results are shown in Figure 6, the gains of in-context demonstrations quickly plateaus when the number of demonstrations just above 3 demonstrations.
-Performance across field types.
-To better understand the task difficulty based on field types, we show the performance of GPT4 (text input; 7 demonstrations) for each input field type in Figure 8.
-We distinguish between "text" and "textarea" fields because they are defined with different HTML tags (<input type="text"> vs.
+4.4 Analysis: Performance Per Field Types
+We provide extended details on performances per input field type in Table 5.
+Note that we distinguish between "text" and "textarea" fields because they are defined with different HTML tags (<input type="text"> vs.
 <textarea>).
 The former is typically used for short inputs, while the latter is for longer, multi-sentence outputs.
-A notable finding was the surprisingly lower performance of GPT4 compared to InternVL2 on “text” input fields, which we could not explain.
+There is no consistent patter here.
+Most models tend to have similar weaknesses and strengths, which is not surprising as they’re generally trained on similar-ish [pre-training] data.
+A notable finding was the surprisingly lower performance of GPT4 compared to other models on “text” input fields, which we could not explain.
 However, since the data is skewed toward checkboxes (Figure 3), GPT4 performs better in the aggregate evaluations.
+This suggests the potential for hybrid models by combining those that excel at specific input field types.
+Model | # of params | # of demos | Modality | all | checkbox | radio | text | textarea |
+Llama-3.2 | 3 B | 3 | T | 20.8 | 58.7 | 25.8 | 19.1 | 0.3 |
+Llama-3.2 | 3 B | 7 | T | 20.3 | 56.8 | 22.9 | 22.0 | 2.9 |
+Gemma-2 | 7 B | 3 | T | 20.4 | 54.1 | 28.7 | 15.1 | 2.0 |
+InternVL | 40 B | 3 | T+ V | 30.7 | 59.0 | 85.0 | 46.0 | 12.0 |
+InternVL | 40 B | 7 | T+ V | 26.1 | 85.0 | 28.2 | 48.1 | 7.5 |
+InternVL | 76 B | 3 | T+ V | 30.4 | 63.7 | 29.7 | 48.4 | 17.1 |
+InternVL | 76 B | 7 | T+ V | 29.6 | 56.6 | 30.5 | 46.4 | 12.3 |
+GPT4-VL | - | 3 | T+ V | 30.3 | 73.3 | 34.2 | 16.4 | 15.5 |
 4.5 Error Analysis
@@
 However, the evaluation focused on the content of the responses rather than their syntactical correctness, so syntax errors did not affect the assessment of answer accuracy.
-We would like to provide more details upon having more space.
+In Table 6 we highlight examples of model predictions for a given UI.
 5 Conclusion and Future Work
@@
 We consider these extensions somewhat orthogonal to the primary focus of this work and hope future research will address them.
-Limitations.
-We discuss several limitations: (i) No navigation between pages: Our benchmark does not require navigation between web pages.
+Limitations
+We discuss several limitations:
+Intra page navigation.
+Our benchmark does not require navigation between web pages.
 While inter-page navigation is important, it is not the only challenge.
 Effective understanding and manipulation of each page, which is our focus, remains a significant challenge for web agents.
-Our benchmark still requires navigation within each page, involving multiple rounds of interaction to solve tasks with multiple input fields.
-As shown in our experiments, this remains challenging for the models.
+Our benchmark still requires navigation within each page.
+This is because almost all of our tasks have more than one step (input field), each needed to be solved in a different round of interaction.
+As shown in our experiments, the benchmark remains challenging for the models.
 We accept this trade-off to obtain more natural tasks compared to most existing benchmarks.
-(ii) Simplified evaluation: This is not an inherent limitation of our benchmark but a simplifying assumption in our evaluation.
+Simplified evaluation.
+This is not an inherent limitation of our benchmark but a simplifying assumption in our evaluation.
 For simplicity, our evaluation setup provides some guidance by specifying the names of the fields to be modified.
 Future work should explore variants of our experiments where such hints are not provided to the model.
-Ethical considerations.
+Scope.
+Our benchmark’s distribution is biased to crowd-sourcing tasks and hence it is not a holistic measure of web-based agents.
+That being said, all benchmarks carry their own biases.
+In the context of web-based tasks, all the existing benchmarks (WebShop, Mind2Web, WebArena, etc.) have their own set of assumptions and restrictions about the scope of web-based agents.
+We view all these efforts to be complementary, each quantifying a unique aspect of intelligent behavior on the web.
+Ethical Considerations
 We recognize concerns that this work could lead to technologies replacing crowd workers, who are vital to AI development.
@@
 GPU machines for conducting experiments are provided by the ARCH (Rockfish) cluster (https://www.arch.jhu.edu).
+Supplemental Material
+Appendix A Additional Related Work
+Here we discuss other related work that did not fit in the main text.
+Multi-modal reasoning tasks.
+TurkingBench is also related to efforts in multi-modal interactive environments (Gur et al., 2018; Ku et al., 2020; Li et al., 2020, 2022; Li and Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021).
+However, these often feature simple instructions that are only a few sentences long, unlike our more extensive instructions embedded within web pages.
+Web-based agents.
+The concept of intelligent automated assistant agents collaborating with humans to complete tasks has been around for some time (Allen et al., 2007) and can be seen as an extension of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022).
+Recent literature has explored various forms of supervision, including behavior cloning of actions (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022; Liu et al., 2023b), and in-context learning (ICL) (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023).
+Our work focuses on introducing a new benchmark, providing ICL baselines, and leaving the exploration of more sophisticated models for future research.
+Appendix B Pseudocode for the oracle baseline
+Appendix C Extracting the “relevant” HTML code for each field
+We detail the procedure for retrieving "relevant" neighboring HTML, as discussed in §4 under "Encoding the tasks for evaluation." The “relevant” HTML encoding is implemented by extracting and returning a subset of HTML adjacent to (above and under) the specific input field on a webpage.Specifically: (1) We retrieve the full HTML content of the target elements’ grandparents elements.
+This potentially contains a lot of content.
+(2.) We split this code into html elements (split based on occurrence of ">") (3) We select 15 elements before and 30 elements below the target element.
+The original function is accessible in our public code base.
+Category: Both automatic and human evaluations classified the action as incorrect (true negative). |
+Explanation: It is perfectly legal to do carry your pet in public, even as a status symbol. |
+Category: Both automatic and human evaluations classified the action as incorrect (true negative). |
+Explanation: Our society does not "strongly pressure" people to carry pets in public. |
+Category: Automatic evaluation rated the action as incorrect but human evaluations rated it correct (false negative). |
+Explanation: The automatic evaluation has rated it incorrect since ‘previous’ was not found in the list of gold answers: [’filled’, ’this’, ’last’]. |
+Category: Automatic evaluation rated the action as incorrect but human evaluations rated it correct (false negative). |
+Explanation: The automatic evaluation has rated it incorrect since ‘great’ was not found in the list of gold answers: [’good’, ’guy’, ’girl’, ’Good’, ’big’]. |
+Category: Automatic evaluation and human evaluation both rated the action as correct (true positive). |
+Explanation: – |
+Category: Automatic evaluation and human evaluation both rated the action as correct (true positive). |
+Explanation: – |
+Category: Automatic evaluation rated the action as correct even though human evaluations rated it incorrect (false positive). |
+Explanation: The statement is true, but the highlighted region is irrelevant. |
+Category: The model response does not adhere to the expected syntax. |
+Explanation: The execution here failed since instead of selecting one of the radio values from -2 to 2, the model responded with a blank string. |